# The Enduring Relevance of Semiempirical Quantum Mechanics

### v2

@@ 3 Hamiltonian model forms
SQM and QM method development have been operating at different activity levels that have resulted in different standard practices. Each popular SQM model family has been developed independently by a relatively small group of researchers and implemented in its own software with its own standards. Only very recently has this situation begun to change with software such as Sparrow 80 and ULYSSES 81 providing implementations of multiple SQM model families in a common code base with a common interface. In contrast, there are many QM software programs that implement either the Gaussian AO or plane-wave pseudopotential formalisms. Within those formalisms, there are many distinct instances of standard components such as Gaussian AO basis sets, pseudopotentials, and DFT exchange-correlation functionals. To accommodate and encourage this level of activity, these components have standard formal structures and often standard data formats, which enables their large-scale aggregation such as in the Basis Set Exchange website 52 or the Libxc software library 55. More standards in SQM model structures could similarly encourage more activity, and enable more interoperability between different SQM development efforts and between QM and SQM efforts. In particular, a common model form for many-electron Hamiltonians could enable QM software to apply the same methods to both QM and SQM Hamiltonians. Although many QM software programs contain implementations of SQM models, they are usually isolated components that are incompatible with most of the other QM features.
Practically, a more standard and interoperable form for both QM and SQM Hamiltonians corresponds to a standard form for the important matrix elements that define these Hamiltonians. Here, I consider the specific case of AO-based matrix elements that are relevant to all popular SQM models and QM software using the Gaussian AO formalism. To enable parameterized models with a manageable number of parameters, the structured form of SQM matrix elements must be designed to minimize the amount of data whereas QM matrix element structure is usually designed to minimize approximation error. In particular, the structured form that I propose here is not exactly compatible with standard factorizations of four-center Coulomb integrals in QM, but it is a systematically improvable approximation that has computational benefits inspired by the efficiency of SQM calculations.
@@ 3.1 Two-center matrix elements
for and . are matrix elements between real spherical harmonics of standard alignment to and aligned to , which can be evaluated using fast, stable recursive formulas 82. are real-valued Slater–Koster model parameters with a symmetry, , that reduces the set of independent parameters to .
The main benefit of using the Slater–Koster formalism is its flexibility. can be evaluated as integrals of AOs in a standard pose with as the distance vector from to , or by interpolation from values in a Slater–Koster table, or from a parameterized SQM formula. Typically, they are truncated beyond a cutoff radius, either to zero for local matrix elements or to a multipole-multipole interaction for Coulomb matrix elements. While this formalism can be used with orthogonal AOs, the use of non-orthogonal AOs with overlap matrix elements is empirically known to improve transferability 83, 84, and it is formally required to maintain and as good quantum numbers. This flexibility also makes it easier to reconsider and transition from old SQM model forms such as the Klopman–Ohno electrostatic interaction, which corresponds to charge distributions with unphysical algebraic tails that significantly complicate periodic electrostatics 85. It also accommodates a broad range of parametric complexity, from general distance dependence for each pair of elements and shells to a single Slater orbital exponent or Klopman–Ohno radius for each element and shell as in the MNDO-family models 31. Having a single length-scale parameter for each element and shell is appropriate for models that are fit mostly to equilibrium structures with a limited variety of nearest-neighbor distances.
@@ 3.2 Four-center matrix elements
It is popular in quantum chemistry to fit a general three-center to minimize errors in a chosen auxiliary AO basis. In SQM models, a general has an intractable number of parameters to fit, and it is more practical to restrict to the same atoms as and . Such two-center density fitting has a long history 86 and modern versions that attain high accuracy 87.
@@
where is the one-center AO overlap matrix, which is typically the identity matrix, and are the total charges of the auxiliary AO basis functions. For explicit function approximations, this fit can be achieved by a local approximation using the Coulomb metric with multipole moment constraints 88.
@@
where is the matrix inverse of the corresponding two-center matrix block of . Explicit function approximations can constrain higher multipole moments to reduce errors 88.
Overall, this SQM model of four-center Coulomb integrals depends only on two-center, two-index parameters and and one-center, three-index parameters and . This factored tensor structure maintains a modest parametric complexity and enables the use of fast solver algorithms based on intermediate tensor contractions 67. It reduces to the NDDO approximation 45 in the case of orthogonal orbitals, and the NDDO approximation is the main reason why SQM calculations are more efficient than QM calculations in a minimal AO basis. The NDDO approximation uses orthogonal orbitals to eliminate interatomic monopole moments, but it is fundamentally incapable of accounting for higher-order interatomic multipole moments. While the NDDO approximation is reasonably accurate for individual Coulomb integrals, large overall errors can accumulate from a large number of individual integral errors, and increasing the number of AOs does not reduce these errors 89. This more general form retains the computational benefits of NDDO while enabling a systematic reduction of errors by expanding the number of primary and auxiliary AOs. Besides the NDDO approximation in the MNDO-family SQM models, most popular SQM models retain only a monopole approximation of the Coulomb interaction except for the GFN2-xTB model 90, which uses a multipole approximation up to quadrupoles with parameterized interatomic multipole moments and structure similar to Eq. (8).
@@ 4 Ongoing work and challenges
The NRL tight-binding model 36 was the most substantial open effort to build SQM models for materials, but it was limited to unary systems for 53 elements of the periodic table. The ThreeBodyTB.jl software 91 is ongoing work to extend the SQM modeling of materials to binary systems comprised of 65 elements. Its expanded and refined SQM model contains three-body Hamiltonian corrections and self-consistent atomic charges to improve transferability, and it is fit to total energies and band structures of DFT reference data.
There is ongoing work to incorporate QM reference data more directly into SQM models through a more explicit Hamiltonian downfolding process into a Natural Orbital Tied Constructed Hamiltonian (NOTCH) form 92. Similar efforts date back to Dewar’s post-AM1 work on the SAM1 model 93, and NOTCH is an increase in technical sophistication that replaces more of the empirical fitting with a formal transformation between basis sets. The NOTCH form includes a complete set of two-center Coulomb integrals instead of using the NDDO approximation, but it still neglects three-center and four-center Coulomb integrals. So far, NOTCH has only been tested on diatomic molecules, which are not sensitive to these neglected Coulomb integrals, and more transferable versions of NOTCH may need to approximate these integrals rather than neglect them.
Other ongoing work in SQM model construction is drawing from a larger set of components such as expanded AO basis sets with split-valence and polarization functions and approximate, fixed-cost SCF solvers 66. Such work highlights the need to better understand sources of error in SQM models to budget costs and errors more effectively with an increasing number of model design choices. The larger basis set increases cost and accuracy while the fixed-cost SCF solver prevents systems with slow SCF convergence from becoming cost outliers, but it may convert these cost outliers into error outliers and further broaden the tails of error distributions. Notably, this more modern SQM model form now has enough control over sources of error to benefit from larger basis sets, unlike older efforts to construct SQM models beyond a minimal basis 94, 95.
There is also ongoing work to refine SQM software in addition to SQM models, particularly within the GFN family of models. The tblite library 96 is a result of work to increase the modularity of xTB, which provides convenient access to GFN model Hamiltonians for other SQM and QM software. The dxtb framework 97 is a Python implementation of GFN models using automatic differentiation to support refitting of model parameters to custom reference data. The SQMBox library 98 tightly integrates GFN models into the TeraChem software through the elementary operation of mapping from a density matrix to a Fock matrix, which may also enable integration with other QM software of a similar design.
@@ 4.1 Role of machine learning
In the first category, the development of established SQM model families is continuing through the addition of new ML corrections to the total energy, often replacing simpler corrections based on physical models. The AIQM1 model 99 continues the development of orthogonalization-corrected MNDO-family models by replacing dispersion corrections with a modified version of the ANI-1x model 100. The PM6-ML model 101 continues the development of the PM model family by replacing hydrogen-bond corrections with a TorchMD-NET potential 102. The QD model 103 corrects a DFTB model using a DeepPot-SE potential 104. In some cases, the combination of SQM and ML models is inspiring new ML model forms such as the ChIMES cluster expansion 105, which has been used as a flexible alternate form for the repulsive correction in DFTB models.
@@
In the second category, there is more diverse research activity because there are many ways to combine ML and SQM models. Static parameters of an SQM model can be replaced by environment-dependent parameters generated by an ML model, as demonstrated by a combination of the PM3 model and a deep neural network 106, which is a further increase in model flexibility relative to older charge-dependent SQM parameters 107. The sophisticated optimization machinery used to fit ML models can be adapted to SQM model forms, such as with the direct optimization of Slater-Koster parameters in a spline representation 108. ML software frameworks can be used to implement existing SQM models and prepare them for combined optimization alongside ML corrections, which is the approach of the TBMaLT software 109. While most of this activity is focused on AO-based models, there is even some renewed interest in EPM models, with the use of equivariant neural networks to generate more transferable EPM model parameters 110.
@@
The research activity in combining ML and SQM models is likely to increase for the foreseeable future, but it should remain mindful of two important caveats. First, the two SQM model families that have achieved the widest coverage of the periodic table, PM 111 and GFN 34, use matrix elements with monoatomic parameters and simple distance dependence in their Hamiltonian model forms. The general form of Slater–Koster matrix elements allows for diatomic parameters and arbitrary distance dependence, but only recently has there been enough QM reference data to fit these more general forms 91. Monoatomic and diatomic parameters correspond to linear and quadratic growth in the number of model parameters with the number of elements covered in the periodic table. In general, ML models are likely to have a combinatorial growth of parameters if they maintain a constant accuracy while increasing elemental coverage, and it will be challenging to generate enough reference data to fit this growing number of parameters. Finally, ML models are often considered to have a negligible computational cost relative to SQM models, but these costs are not always well separated and might eventually cross as ML models become more complicated. For example, PM6-ML calculations of large proteins using the fast MOZYME solver spend 13% of their total computational cost on evaluating their ML correction 101. This is part of a general trend in MM, where the cost of interatomic potentials has been steadily growing as their complexity and accuracy have both increased 112. If the cost of the ML component exceeds the cost of the SQM component in a hybrid model, then a better cost-accuracy balance might be achieved by switching to a more accurate and expensive SQM component.
@@ 4.2 Hydrogen cluster example
Before concluding this paper, I consider a simple numerical example that shows the value of understanding errors in individual SQM model components. Hydrogen is the most prevalent element in biomolecular simulations, so the cost and accuracy of SQM models in biomolecular applications is sensitive to how hydrogen is modeled. While most AO-based SQM models consider only one orbital on each hydrogen atom, there are models that add a second orbital 34, or a shell of orbitals 113, or both 66. However, it is difficult to separate the effect of that specific design decision from all of the other differences in model design and fitting. Here, I examine some of the limits of an SQM model of hydrogen in a minimal AO basis.
I make several simplifying assumptions in this example. First, I consider a model Hamiltonian using Slater–Koster parameters and factored four-center Coulomb integrals from Eq. (5) with a single -type primary and auxiliary AO on each hydrogen atom. This minimal basis simplifies the density fitting structure, reducing the model to two one-center parameters—orbital energy and electronic Coulomb self-energy—and five two-center parameters—overlap (), one-body Hamiltonian (), and electron-electron (), electron-proton (), and proton-proton () Coulomb interactions. I use the exact AO energy for an isolated atom, leaving one one-center parameter and five two-center parameters. Second, I model electron correlation exactly using full configuration interaction to isolate the Hamiltonian model errors. Third, I restrict the reference data to the hydrogen dimer and the symmetric hydrogen trimer so that the Slater–Koster parameters can be fit independently for each interatomic distance. I then generate FCI/def2-QZVPP reference data using PySCF 47 for all stationary electronic states in the single-occupancy sector of the energy spectrum corresponding to the number of electronic configurations in a minimal AO basis with up to one electron per atom.
@@ 5 Conclusions
As evident in the collapsing market share of SQM software over the last three decades, development and application of atomistic simulation has substantially polarized towards the extremes of high-accuracy QM and low-cost MM. The rapidly growing popularity of ML and quantum computing research might be interpreted as a further polarization, with QM methods now using expensive quantum computing hardware in the pursuit of higher accuracy 114 and MM methods now replacing even more physical modeling with data-driven ML models. Atomistic method development would perhaps be better served by an interconnected pipeline from high-cost, high-accuracy QM methods to SQM models of lower accuracy or narrower applicability before fragmenting into a constellation of special-purpose MM models. There have already been multiple efforts over the last three decades to encourage and advocate for more SQM research 115, 116, 117, 118, 119, 29 that have not arrested this collapse. I have argued in this paper for the enduring relevance of SQM to atomistic simulation, regardless of its decline.
Even without a substantial increase in academic activity, the steadily growing market for atomistic simulation might eventually be large enough to support SQM method and software development solely through commercial demand. While SQM software presently has a very small market share, latent demand is difficult to quantify, and new SQM software that is more responsive to demand might eventually increase its market share once again. Historically, SQM models have mostly been products of academic research that were freely distributed, but some SQM models such as the QUASINANO parameter sets 120, 121 are now also being developed for commercial distribution. A financially sustainable future for SQM method and software development may necessitate a careful partitioning between open and closed components.

### v3

@@ 3 Hamiltonian model forms
SQM and QM method development have been operating at different activity levels that have resulted in different standard practices. Each popular SQM model family has been developed independently by a relatively small group of researchers and implemented in its own software with its own standards. Only very recently has this situation begun to change with software such as Sparrow 78 and ULYSSES 79 providing implementations of multiple SQM model families in a common code base with a common interface. In contrast, there are many QM software programs that implement either the Gaussian AO or plane-wave pseudopotential formalisms. Within those formalisms, there are many distinct instances of standard components such as Gaussian AO basis sets, pseudopotentials, and DFT exchange-correlation functionals. To accommodate and encourage this level of activity, these components have standard formal structures and often standard data formats, which enables their large-scale aggregation such as in the Basis Set Exchange website 51 or the Libxc software library 54. More standards in SQM model structures could similarly encourage more activity, and enable more interoperability between different SQM development efforts and between QM and SQM efforts. In particular, a common model form for many-electron Hamiltonians could enable QM software to apply the same methods to both QM and SQM Hamiltonians. Although many QM software programs contain implementations of SQM models, they are usually isolated components that are incompatible with most of the other QM features.
Practically, a more standard and interoperable form for both QM and SQM Hamiltonians corresponds to a standard form for the important matrix elements that define these Hamiltonians. Here, I consider the specific case of AO-based matrix elements that are relevant to all popular SQM models and all QM software using the Gaussian AO formalism. To enable parameterized models with a manageable number of parameters, the structured form of SQM matrix elements must be designed to minimize the amount of data whereas QM matrix element structure is usually designed to minimize approximation error. In particular, the structured form that I propose here is not exactly compatible with standard factorizations of four-center Coulomb integrals in QM, but it is a systematically improvable approximation that has computational benefits inspired by the efficiency of SQM calculations.
@@ 3.1 Two-center matrix elements
for and . are matrix elements between real spherical harmonics of standard alignment to and aligned to , which can be evaluated using fast, stable recursive formulas 80. are real-valued Slater–Koster model parameters with a symmetry, , that reduces the set of independent parameters to .
The main benefit of using the Slater–Koster formalism is its flexibility. can be evaluated as integrals of AOs in a standard pose with as the distance vector from to , or by interpolation from values in a Slater–Koster table, or from a parameterized SQM formula. Typically, they are truncated beyond a cutoff radius, either to zero for local matrix elements or to a multipole-multipole interaction for Coulomb matrix elements. While this formalism can be used with orthogonal AOs, the use of non-orthogonal AOs with overlap matrix elements is empirically known to improve transferability 81, 82, and it is formally required to maintain and as good quantum numbers. This flexibility also makes it easier to reconsider and transition from old SQM model forms such as the Klopman–Ohno electrostatic interaction, which corresponds to charge distributions with unphysical algebraic tails that significantly complicate periodic electrostatics 83. It also accommodates a broad range of parametric complexity, from general distance dependence for each pair of elements and shells to a single Slater orbital exponent or Klopman–Ohno radius for each element and shell as in the MNDO-family models 6. Having a single length-scale parameter for each element and shell is appropriate for models that are fit mostly to equilibrium structures with a limited variety of nearest-neighbor distances.
@@ 3.2 Four-center matrix elements
It is popular in quantum chemistry to fit a general three-center to minimize errors in a chosen auxiliary AO basis. In SQM models, a general has an intractable number of parameters to fit, and it is more practical to restrict to the same atoms as and . Such two-center density fitting has a long history 84 and modern versions that attain high accuracy 85.
@@
where is the one-center AO overlap matrix, which is typically the identity matrix, and are the total charges of the auxiliary AO basis functions. For explicit function approximations, this fit can be achieved by a local approximation using the Coulomb metric with multipole moment constraints 86.
@@
where is the matrix inverse of the corresponding two-center matrix block of . Explicit function approximations can constrain higher multipole moments to reduce errors 86.
Overall, this SQM model of four-center Coulomb integrals depends only on two-center, two-index parameters and and one-center, three-index parameters and . This factored tensor structure maintains a modest parametric complexity and enables the use of fast solver algorithms based on intermediate tensor contractions 66. It reduces to the NDDO approximation 44 in the case of orthogonal orbitals, and the NDDO approximation is the main reason why SQM calculations are more efficient than QM calculations in a minimal AO basis. The NDDO approximation uses orthogonal orbitals to eliminate interatomic monopole moments, but it is fundamentally incapable of accounting for higher-order interatomic multipole moments. While the NDDO approximation is reasonably accurate for individual Coulomb integrals, large overall errors can accumulate from a large number of individual integral errors, and increasing the number of AOs does not reduce these errors 87. This more general form retains the computational benefits of NDDO while enabling a systematic reduction of errors by expanding the number of primary and auxiliary AOs. Besides the NDDO approximation in the MNDO-family SQM models, most popular SQM models retain only a monopole approximation of the Coulomb interaction except for the GFN2-xTB model 88, which uses a multipole approximation up to quadrupoles with parameterized interatomic multipole moments and structure similar to Eq. (8).
@@ 4 Ongoing work and challenges
The NRL tight-binding model 35 was the most substantial open effort to build SQM models for materials, but it was limited to unary systems for 53 elements of the periodic table. The ThreeBodyTB.jl software 89 is ongoing work to extend the SQM modeling of materials to binary systems comprised of 65 elements. Its expanded and refined SQM model contains three-body Hamiltonian corrections and self-consistent atomic charges to improve transferability, and it is fit to total energies and band structures of DFT reference data.
There is ongoing work to incorporate QM reference data more directly into SQM models through a more explicit Hamiltonian downfolding process into a Natural Orbital Tied Constructed Hamiltonian (NOTCH) form 90. Similar efforts date back to Dewar’s post-AM1 work on the SAM1 model 91, and NOTCH is an increase in technical sophistication that replaces more of the empirical fitting with a formal transformation between basis sets. The NOTCH form includes a complete set of two-center Coulomb integrals instead of using the NDDO approximation, but it still neglects three-center and four-center Coulomb integrals. So far, NOTCH has only been tested on diatomic molecules, which are not sensitive to these neglected Coulomb integrals, and more transferable versions of NOTCH may need to approximate these integrals rather than neglect them.
Other ongoing work in SQM model construction is drawing from a larger set of components such as expanded AO basis sets with split-valence and polarization functions and approximate, fixed-cost SCF solvers 65. Such work highlights the need to better understand sources of error in SQM models to budget costs and errors more effectively with an increasing number of model design choices. The larger basis set increases cost and accuracy while the fixed-cost SCF solver prevents systems with slow SCF convergence from becoming cost outliers, but it may convert these cost outliers into error outliers and further broaden the tails of error distributions. Notably, this more modern SQM model form now has enough control over sources of error to benefit from larger basis sets, unlike older efforts to construct SQM models beyond a minimal basis 92, 93.
There is also ongoing work to refine SQM software in addition to SQM models, particularly within the GFN family of models. The tblite library 94 is a result of work to increase the modularity of xTB, which provides convenient access to GFN model Hamiltonians for other SQM and QM software. The dxtb framework 95 is a Python implementation of GFN models using automatic differentiation to support refitting of model parameters to custom reference data. The SQMBox library 96 tightly integrates GFN models into the TeraChem software through the elementary operation of mapping from a density matrix to a Fock matrix, which may also enable integration with other QM software of a similar design.
@@ 4.1 Role of machine learning
In the first category, the development of established SQM model families is continuing through the addition of new ML corrections to the total energy, often replacing simpler corrections based on physical models. The AIQM1 model 97 continues the development of orthogonalization-corrected MNDO-family models by replacing dispersion corrections with a modified version of the ANI-1x model 98. The PM6-ML model 99 continues the development of the PM model family by replacing hydrogen-bond corrections with a TorchMD-NET potential 100. The QD model 101 corrects a DFTB model using a DeepPot-SE potential 102. In some cases, the combination of SQM and ML models is inspiring new ML model forms such as the ChIMES cluster expansion 103, which has been used as a flexible alternate form for the repulsive correction in DFTB models.
@@
In the second category, there is more diverse research activity because there are many ways to combine ML and SQM models. Static parameters of an SQM model can be replaced by environment-dependent parameters generated by an ML model, as demonstrated by a combination of the PM3 model and a deep neural network 104, which is a further increase in model flexibility relative to older charge-dependent SQM parameters 105. The sophisticated optimization machinery used to fit ML models can be adapted to SQM model forms, such as with the direct optimization of Slater-Koster parameters in a spline representation 106. ML software frameworks can be used to implement existing SQM models and prepare them for combined optimization alongside ML corrections, which is the approach of the TBMaLT software 107. While most of this activity is focused on AO-based models, there is even some renewed interest in EPM models, with the use of equivariant neural networks to generate more transferable EPM model parameters 108.
@@
The research activity in combining ML and SQM models is likely to increase for the foreseeable future, but it should remain mindful of two important caveats. First, the two SQM model families that have achieved the widest coverage of the periodic table, PM 109 and GFN 33, use matrix elements with monoatomic parameters and simple distance dependence in their Hamiltonian model forms. The general form of Slater–Koster matrix elements allows for diatomic parameters and arbitrary distance dependence, but only recently has there been enough QM reference data to fit these more general forms 89. Monoatomic and diatomic parameters correspond to linear and quadratic growth in the number of model parameters with the number of elements covered in the periodic table. In general, ML models are likely to have a combinatorial growth of parameters if they maintain a constant accuracy while increasing elemental coverage, and it will be challenging to generate enough reference data to fit this growing number of parameters. Finally, ML models are often considered to have a negligible computational cost relative to SQM models, but these costs are not always well separated and might eventually cross as ML models become more complicated. For example, PM6-ML calculations of large proteins using the fast MOZYME solver spend 13% of their total computational cost on evaluating their ML correction 99. This is part of a general trend in MM, where the cost of interatomic potentials has been steadily growing as their complexity and accuracy have both increased 110. If the cost of the ML component exceeds the cost of the SQM component in a hybrid model, then a better cost-accuracy balance might be achieved by switching to a more accurate and expensive SQM component.
@@ 4.2 Hydrogen cluster example
Before concluding this paper, I consider a simple numerical example that shows the value of understanding errors in individual SQM model components. Hydrogen is the most prevalent element in biomolecular simulations, so the cost and accuracy of SQM models in biomolecular applications is sensitive to how hydrogen is modeled. While most AO-based SQM models consider only one orbital on each hydrogen atom, there are models that add a second orbital 33, or a shell of orbitals 111, or both 65. However, it is difficult to separate the effect of that specific design decision from all of the other differences in model design and fitting. Here, I examine some of the limits of an SQM model of hydrogen with a minimal AO basis.
I make several simplifying assumptions in this example. First, I consider a model Hamiltonian using Slater–Koster parameters and factored four-center Coulomb integrals from Eq. (5) with a single -type primary and auxiliary AO on each hydrogen atom. This minimal basis simplifies the density fitting structure, reducing the model to two one-center parameters—orbital energy and electronic Coulomb self-energy—and five two-center parameters—overlap (), one-body Hamiltonian (), and electron-electron (), electron-proton (), and proton-proton () Coulomb interactions. I use the exact AO energy for an isolated atom, leaving one one-center parameter and five two-center parameters. Second, I model electron correlation exactly using full configuration interaction to isolate the Hamiltonian model errors. Third, I restrict the reference data to the hydrogen dimer and the symmetric hydrogen trimer so that the Slater–Koster parameters can be fit independently for each interatomic distance. I then generate FCI/def2-QZVPP reference data using PySCF 46 for all stationary electronic states in the single-occupancy sector of the energy spectrum corresponding to the number of electronic configurations in a minimal AO basis with up to one electron per atom.
@@ 5 Conclusions
As evident in the collapsing market share of SQM software over the last three decades, development and application of atomistic simulation has substantially polarized towards the extremes of high-accuracy QM and low-cost MM. The rapidly growing popularity of ML and quantum computing research might be interpreted as a further polarization, with QM methods now using expensive quantum computing hardware in the pursuit of higher accuracy 112 and MM methods now replacing even more physical modeling with data-driven ML models. Atomistic method development would perhaps be better served by an interconnected pipeline from high-cost, high-accuracy QM methods to SQM models of lower accuracy or narrower applicability before fragmenting into a constellation of special-purpose MM models. There have already been multiple efforts over the last three decades to encourage and advocate for more SQM research 113, 114, 115, 116, 117, 29 that have not arrested this collapse. I have argued in this paper for the enduring relevance of SQM to atomistic simulation, regardless of its decline.
Even without a substantial increase in academic activity, the steadily growing market for atomistic simulation might eventually be large enough to support SQM method and software development solely through commercial demand. While SQM software presently has a very small market share, latent demand is difficult to quantify, and new SQM software that is more responsive to demand might eventually increase its market share once again. Historically, SQM models have mostly been products of academic research that were freely distributed, but some SQM models such as the QUASINANO parameter sets 118, 119 are now also being developed for commercial distribution. A financially sustainable future for SQM method and software development may necessitate a careful partitioning between open and closed components.
6 Biography
Jonathan E. Moussa received his doctoral degree in Physics from the University of California, Berkeley in 2008 under the supervision of Prof. Marvin Cohen. His doctoral work focused on applying density functional theory to predict new phonon-mediated superconductors. After two postdoctoral positions in computational materials science, at the University of Texas, Austin with Prof. James Chelikowsky and at Sandia National Laboratories with Dr. Peter Schultz, he switched fields to quantum information science and converted to a staff scientist at the Labs in 2014. In 2018, he joined the Molecular Sciences Software Institute as a software scientist and returned to computational materials science and chemistry. In this role, he led the transition of MOPAC from commercial to open-source software and its ongoing development, maintenance, and support. His main research interest is the balance between cost and accuracy in algorithms and software for atomistic simulation, particularly at the interface between quantum and classical in methods and computers.